A researcher studying the nutritional value of a new candy places a 6.20 g

6.20 g sample of the candy inside a bomb calorimeter and combusts it in excess oxygen. The observed temperature increase is 2.06 ∘C. 2.06 ∘C. If the heat capacity of the calorimeter is 36.70 kJ⋅K−1, 36.70 kJ⋅K−1, how many nutritional Calories are there per gram of the candy?


1
Expert's answer
2021-04-07T11:32:19-0400

Heat gained by the calorimeter is


"Q=\varDelta t*C=2.06 ^o C *36.7\dfrac{kJ}{K}=75.6 kJ"


1 cal is 4.184 J, so


"75.6kJ=18069 cal"


"\dfrac{18069cal}{6.20g}=2914.4\dfrac{cal}{g}"